OpenStudy (anonymous):

Find the product and simply -2x/(3y^2) × 2y/(5)

OpenStudy (anonymous):

Would it be -4x/(15y) ????

OpenStudy (anonymous):

is it \[\huge \frac{-2x}{3y^2}\times \frac{2y}{5}\]

OpenStudy (anonymous):

if so, then your answer is right

OpenStudy (anonymous):

So -4x/ (15y) is correct ?